one number is 4 more than twice the second number. Their total is 25. Find the number

Answer by checkley71(8403)

x+y=25
x=2y+4
2y+4+y=25
3y=25-4
3y=21
y=21/3
y=7 answer.
x+7=25
x=25-7
x=18 answer.